OpenStudy (anonymous):

how do you derive the derivative of cotx? I know how to do sin, cos, csc, sec, tan

OpenStudy (anonymous):

d/dx (cotx) = -csc^2(x)

OpenStudy (anonymous):

how do you derive it? I know what it is.

OpenStudy (bahrom7893):

use u sub: u = Sinx du = Cosx

OpenStudy (anonymous):

oh. My mistake. Well, one way would be to use the quotient rule. So cot(x) = cos(x)/sin(x). Taking the derivative of cos(x)/sin(x) using the quotient rule yields: [sin(x)[d/dx(cos(x)] - cos(x)[d/dx(sin(x))]]/sin^2(x)

OpenStudy (bahrom7893):

rewrite Cot = Cos/Sin

OpenStudy (bahrom7893):

please fan us both if we helped!

OpenStudy (anonymous):

What identity does it become? Thats where i got confused in the quotient rule

OpenStudy (bahrom7893):

Okay it becomes: Integral Cot = Integral Cos/Sin Let u = sin du = Cos dx Integral Cot = Integral du/u = Ln|u| + C = Ln|Sin(x)|+C

OpenStudy (anonymous):

its calc 1, I havent done integrals yet

OpenStudy (bahrom7893):

WOOPS LOL sorry i misread the question

OpenStudy (bahrom7893):

Use Cot = Cos*(Sin)^(-1) And use product rule

OpenStudy (anonymous):

from where I left off: the top becomes -sin^2(x) - cos^2(x) = -1 then the whole thing becomes -1/sin^2(x) = -csc^2(x)

OpenStudy (bahrom7893):

Or Cot = Cos/Sin and quotient rule

OpenStudy (anonymous):

beautiful, thanks pyeh9

OpenStudy (anonymous):

my professor wanted us to derive all identities so I was confused on that one for the exam tomorrow

OpenStudy (anonymous):

sure thing!